Customer expectations on pizza delivery are such that most people are happy with a 20 minute delivery and very upset if it takes more than 40 minutes. You have observed the delivery time for a random sample of customers for two days as follows:

Day        1              2              3              4              5              6              7              8              9              10

1              12           23           22           28           34           10           55           20           37           41

2              32           23           27           35           12           48           40           21           35           30

Answer the following questions pertaining to this data (hint: irst calculate the mean and thee standard deviation of this data)

  1. Competition is really cut-throat and in an effort to improve service delivery quality, one proposal calls for disciplining and/or reprimanding the delivery person(s) who takes more than 40 minutes for delivery. What do you think about the proposal?
  2. What is the appropriate measure of process capability here: Cp or Cpk?/ Why?
  3. What is the capability of the process? Is a control chart useful for this situation? Why or why not?
  4. Suggest the upper and lower control limits of a control chart that would monitor the pizza delivery process?
  5. How would you use this control chart?
  6. Suggest at least three process/quality improvement efforts that are appropriate given the above data and observations.
Answer #1
  1. Mean & Standard Deviation

Customer

Day 1

Day 2

Mean & Std. Dev.

1

12

32

2

23

23

3

22

27

4

28

35

5

34

12

6

10

48

7

55

40

8

20

21

9

37

35

10

41

30

Mean

28.20

30.30

29.25

Std. Dev.

13.82

10.22

2.55

Excel Formula Sheet:

Customer

Day 1

Day 2

Mean & Std. Dev.

1

12

32

2

23

23

3

22

27

4

28

35

5

34

12

6

10

48

7

55

40

8

20

21

9

37

35

10

41

30

Mean

=AVERAGE(B2:B11)

=AVERAGE(C2:C11)

=AVERAGE(B12:C12)

Std. Dev.

=STDEV(B2:B11)

=STDEV(C2:C11)

=STDEV(B13:C13)

  1. Competition is really cut-throat and in an effort to improve service delivery quality, one proposal calls for disciplining and/or reprimanding the delivery person(s) who takes more than 40 minutes for delivery. What do you think about the proposal?

I accept, in any business delivery on time and customer satisfaction plays a key role in sales and profit. One who takes more than 40 min can never satisfy a customer, so reprimanding the delivery person(s) is advisable.

  1. What is the appropriate measure of process capability here: Cp or Cpk?/ Why?

Customer

Day 1

Day 2

Mean & Std. Dev.

1

12

32

2

23

23

3

22

27

4

28

35

5

34

12

6

10

48

7

55

40

8

20

21

9

37

35

10

41

30

Mean

28.20

30.30

29.25

Std. Dev.

13.82

10.22

2.55

Upper Limit

55

48

Lower Limit

10

12

Potential Cp

0.543

0.587

Actual Cpk Index

0.439

0.597

Excel Formula sheet:

Customer

Day 1

Day 2

Mean & Std. Dev.

1

12

32

2

23

23

3

22

27

4

28

35

5

34

12

6

10

48

7

55

40

8

20

21

9

37

35

10

41

30

Mean

=AVERAGE(B2:B11)

=AVERAGE(C2:C11)

=AVERAGE(B12:C12)

Std. Dev.

=STDEV(B2:B11)

=STDEV(C2:C11)

=STDEV(B13:C13)

Upper Limit

=MAX(B2:B11)

=MAX(C2:C11)

Lower Limit

=MIN(B2:B11)

=MIN(C2:C11)

Potential Cp

=(B14-B15)/(6*B13)

=(C14-C15)/(6*C13)

Actual Cpk Index

=ABS(MIN((B14-B12)/(3*B13), (B15-B12)/(3*B13)))

=ABS(MIN((C14-C12)/(3*C13), (C15-C12)/(3*C13)))

  1. What is the capability of the process? Is a control chart useful for this situation? Why or why not?

The capability of the process is less than its actual potential: Cp < Cpk

  1. Suggest the upper and lower control limits of a control chart that would monitor the pizza delivery process?

LCL

Delivery Time

UCL

Day 1-1

20

12

40

Day 1-2

20

23

40

Day 1-3

20

22

40

Day 1-4

20

28

40

Day 1-5

20

34

40

Day 1-6

20

10

40

Day 1-7

20

55

40

Day 1-8

20

20

40

Day 1-9

20

37

40

Day 1-10

20

41

40

Day 2-1

20

32

40

Day 2-2

20

23

40

Day 2-3

20

27

40

Day 2-4

20

35

40

Day 2-5

20

12

40

Day 2-6

20

48

40

Day 2-7

20

40

40

Day 2-8

20

21

40

Day 2-9

20

35

40

Day 2-10

20

30

40

  1. How would you use this control chart?

I would enumerate how many delivery took more than 40 min, how many delivery took less than 20 min and how many delivery took time between 20-40 min.

  1. Suggest at least three process/quality improvement efforts that are appropriate given the above data and observations.
  • Process standardization.
  • Pre-prepared pizzas for faster processing and delivery.
  • FIFO technique.